how long does it take to quadruple the investment compound interest of 7.25%

Answer by Boreal(15235)

The rule of 70 says doubling time in years is 70/interest rate, so here the doubling time is not quite 10 years and the quadrupling time twice that or somewhere between 19 and 20
We don't have the number of times it is compounded. Here it is for continuously compounded.
P/Po=e^rt=4, since it is quadrupled
ln both side
rt=ln 4
t=ln 4/0.0725=19.12 years